Consider a step-index fiber for which n1 = 1.475, n2 = 1.460 and a = 25µm. What is the maximum value of θ for which the ray ill be guided by the fiber. 


1
Expert's answer
2021-05-05T17:38:05-0400

Given,


"n_1=1.475\\n_2=1.460"


"a=25 \ \mu m"


Numerical aperture (NA) =


"\sqrt{n_1^2-n_2^2}"


"=\sqrt{(1.475)^2-(1.460)^2}"


"=\sqrt{2.25-2.19}=0.210"


therefore,


NA= 0.210


now,

we calculate "\theta_{max}"


"\theta_{max}=sin^{-1}(0.210)=12.12\degree"


Answer=12.120